### v1

@@ Abstract
We consider the problem where an agent aims to combine the views and insights of different experts’ models. Specifically, each expert proposes a diffusion process over a finite time horizon. The agent then combines the experts’ models by minimising the weighted Kullback-Leibler divergence to each of the experts’ models. We show existence and uniqueness of the barycentre model and proof an explicit representation of the Radon-Nikodym derivative relative to the average drift model. We further allow the agent to include their own constraints, which results in an optimal model that can be seen as a distortion of the experts’ barycentre model to incorporate the agent’s constraints.
Two deep learning algorithms are proposed to find the optimal drift of the combined model, allowing for efficient simulations. The first algorithm aims at learning the optimal drift by matching the change of measure, whereas the second algorithm leverages the notion of elicitability to directly estimate the value function. The paper concludes with a extended application to combine implied volatility smiles models that were estimated on different datasets.
@@
Mixtures of Experts, Kullback-Leibler, Barycentre, Model Combination, Ensemble Model, Deep learning, Elicitability, Volatility Smiles
@@ 1 Introduction
In many situations a modeller or agent aims to find a single model to e.g., predict, forecast, classify, or approximate observed patterns. While model selection procedures provide rankings of the different models’ performances – thus allowing the agent to select one model – an alternative to disregarding all but one model is model combination. In the statistical literature, so-called forecast ensemble or superensemble models have been widely applied to weather prediction as they are known to improve deterministic forecasts, see e.g., [12] and reference therein. A variant of a statistical superensemble that allows for predictive distribution functions is to regress a quantity of interest (e.g., air pressure, temperature) onto forecasts stemming from different models [12]. Averaging different predictions have also been used in machine learning, where early examples include bagging predictors [5], Adaptive, Reweighting and Combining (ARCing) algorithms and Adaboost [6]. To improve the architecture of neural networks, [13] propose neural networks ensembles – averaging different neural network architectures –, which have been applied in e.g., credit scoring [20] and oil price forecasting [21].
This work focuses on model combinations of dynamic stochastic models in continuous-time. A related stream of literature, though in discrete-time, is called dynamic time warping (see, e.g., [3]), which aims at matching or detecting similarities in time series data by finding a “warping path” that minimises a distance between the time series. Applications of dynamic time warping are widespread and range from classification of genome signals [19], to speech recognition [17], and clustering of financial stocks [10].
@@
In this manuscript, we consider a finite number of experts, each having their own belief about the dynamics of a continuous-time diffusion process. The experts’ models are characterised by different probability measures, under which the process follows the experts’ dynamics. An agent then aims at combining the different experts’ models by finding the probability measure that minimises the weighted Kullback-Leibler (KL) divergence to all models. The resulting dynamics of the stochastic process is described by the barycentre model. Moreover, we allow the agent to include their own views, described by expectations of functions of the terminal values of the process or expected running cost, which results in a constrained barycentre model. Key contributions of this work are as follows. First, we derive the optimal value function of our optimisation problem and a succinct representation of the Radon-Nikodym (RN) derivative of the optimal probability measure to, what we term, the average drift measure. We find that the optimal RN derivative is of Escher type albeit different to the classical solution of maximum entropy, see e.g., [9]. Second, we prove that our constraint optimisation problem is equivalent to an optimisation problem of distorting the expert’s barycentre model to include the agent constraints. The latter optimisation problem, that is distorting stochastic processes to include constraints has been studied in [18] and [15]. Third, we propose two deep learning algorithms to approximate the optimal probability measure and thus the dynamics of the constrained barycentre process, allowing for simulation under the constrained barycentre model. Forth, we apply our framework to combining implied volatility smile models that were estimated on different sets of real data.
@@
Here, we utilise the KL divergence to quantify divergences between probability measures on the path space of stochastic processes. Alternatives could be distances on the space of probability measures stemming from optimal transport theory, such as the Wasserstein distance. There are however two caveats: first, calculating a Wasserstein barycentre between random variables is in general an NP hard problem [1], highlighting the curse of dimensionality associated with the Wasserstein distance; second, as we work in continuous-time and over a finite time horizon, the Wasserstein distance need be replaced by the adapted (also called bicausal) Wasserstein distance. The adapted Wasserstein distance between stochastic processes is only known in special cases, indicatively see [4] and [2], while with the KL divergence we obtain a closed-form solution of the optimal RN derivative for arbitrary dimension of the stochastic process.
@@
The manuscript is structure as follows. Section 2 introduces the experts’ models and the agent’s optimisation problem (P). Its solution is presented in Section 2.2, where we additionally characterise the solution to the pure barycentre problem. Key results are the concise representation of the optimal change of measure (Proposition 2.9) and the recasting of optimisation problem (P) into an optimisation problem of minimally distorting the barycentre measure to include the agent’s constraints (Proposition 2.11). In Section 3, we propose two deep learning algorithms to approximate the drift of the optimally combined model. Section 4.1 illustrates and compares these algorithms on simulated examples. A financial application to combining implied volatility smiles models, that were estimated on different datasets, is provided in Section 4.2.
2 Constrained Kullback-Leibler barycentre
@@
This section first introduces the experts’ models and the optimisation problem that the agent aims to tackle. Second, using a stochastic control approach, we solve the optimisation problem and characterise the optimal model. Third, we relate the optimisation problem to a modified barycentre model, in the spirit of [15].
@@ 2.1 Experts’ opinions and agent’s problem
We work on a complete and filtered probability space, denoted by $\left(\Omega,\mathbb{P},{\mathcal{F}},({\mathcal{F}}_{t})_{t\in{[0,T]}}\right)$, on which we have a $d$-dimensional Brownian motion $B=(B_{t})_{t\in{[0,T]}}$ (with independent components) and the filtration is the natural one generated by $B$. We further have an $d$-dimensional process $X=(X_{t})_{t\in{[0,T]}}$ and a set of experts ${\mathcal{K}}=\{1,\dots,K\}$. Each expert $k\in{\mathcal{K}}$ has a probability measure ${\mathbb{P}}^{(k)}$ – which we call the “model” of expert $k$ – that is equivalent to ${\mathbb{P}}$ (${\mathbb{P}}\sim{\mathbb{P}}^{(k)}$), and believes that under ${\mathbb{P}}^{(k)}$ the process $X$ satisfies the SDE
@@
where $W^{(k)}=(W^{(k)}_{t})_{t\in{[0,T]}}$ denotes a $d$-dimensional ${\mathbb{P}}^{(k)}$-Brownian motion, $\mu^{(k)}(t,x):{\mathds{R}}_{+}\times{\mathds{R}}^{d}\to{\mathds{R}}^{d}$, and $\sigma(t,x):{\mathds{R}}_{+}\times{\mathds{R}}^{d}\to{\rm{S}}^{d}_{++}$, where ${\rm{S}}^{d}_{++}$ is the set of $d$-dimensional positive definite matrices. Note that despite requiring the probability measures to be equivalent, we may allow for differing volatilities by introducing a stochastic volatility process that, under each expert’s model, mean-reverts to the volatility model of the expert.
The next assumption guarantees that a strong solution to (1) exists.
@@
The functions $\mu^{(k)}(\cdot,\cdot)$, for all $k\in{\mathcal{K}}$, and $\sigma(\cdot,\cdot)$ satisfy the linear growth and Lipschitz continuity conditions. That is for all ${t\in[0,T]}$ and all $x\in{\mathds{R}}^{d}$, there exists a $C_{1}<\infty$ such that
@@
where ||$\cdot$|| ⋅ | denotes the Euclidean norm and |||||$\cdot$||| | ⋅ | | the Frobenius norm. Moreover, for all ${t\in[0,T]}$ and all $x,y\in{\mathds{R}}^{d}$, there exists a $C_{2}<\infty$ such that
@@
While the experts have differing views – in particular they disagree on the drift of $X$ – an agent wishes to combine the opinions of these experts and assigns each expert a weight $\pi=(\pi_{1},\dots,\pi_{K})$ with $\pi_{k}\in[0,1]$, $k\in{\mathcal{K}}$, and $\sum_{k\in{\mathcal{K}}}\pi_{k}=1$. That is, the agent aims at reflecting the experts’ options and incorporating them with their own belief into a combined probability measure ${\mathbb{Q}}$, which we call the combined model. Specifically, for functions $g\colon{\mathds{R}}_{+}\times{\mathds{R}}^{d}\to{\mathds{R}}$ and $f\colon{\mathds{R}}^{d}\to{\mathds{R}}$ they wish to ensure that ${\mathbb{E}}^{\mathbb{Q}}[f(X_{T})]=0$ and ${\mathbb{E}}^{\mathbb{Q}}[\int_{0}^{T}g(u,X_{u}){\rm d}u]=0$, where we use the notation ${\mathbb{E}}^{{\mathbb{Q}}}[\cdot]$ and ${\mathbb{E}}^{{\mathbb{P}}^{(k)}}[\cdot]$ to denote expectations under the combined and the $k$-th expert’s model, respectively. If $X_{T}$ is univariate and continuously distributed, the choice $f(x)={\mathds{1}}_{\{x<q\}}-\alpha$, $\alpha\in(0,1)$, for example, corresponds to the agent requiring that the Value-at-Risk (VaR) at level $\alpha$ of the combined model equals to $q$. The weights $\pi$ could be, e.g., obtain by computing the posterior probability that the data the agent is using stems from expert-$k$’s model.
We propose that the agent combines the models by finding the (weighted) Kullback-Leibler (KL) barycentre of the expert models subject to the agent’s constraints, thus aims to solve the optimisation problem
@@
denotes the KL divergence of ${\mathbb{Q}}$ relative to ${\mathbb{P}}^{(k)}$, and ${\mathcal{Q}}$ is the set of probability measures parametrised as
@@
${\mathcal{Q}}:=\Bigg{\{}\;{\mathbb{Q}}[\theta]~{}\Big{|}~{}$ $\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{P}}^{(k)}}=\exp\left\{-\tfrac{1}{2}\int_{0}^{T}|\lambda^{(k)}(t,X_{t})|^{2}{\rm d}t-\int_{0}^{T}\lambda^{(k)}(t,X_{t})\,{\rm d}W_{t}^{(k)}\right\}\,,$
$\text{where}\quad\lambda^{(k)}(t,x):=\sigma^{-1}(t,x)\big{(}\mu^{(k)}(t,x)-\theta(t,x)\big{)}\,,$
$\text{and}\quad{\mathbb{E}}^{{\mathbb{P}}^{(k)}}\left[\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{P}}^{(k)}}\right]=1\ \Bigg{\}}\,.$
As ${\mathbb{P}}^{(k)}\sim{\mathbb{P}}$ this set does not depend on $k$. We parameterise a probability measure ${\mathbb{Q}}[\theta]\in{\mathcal{Q}}$ by its drift $\theta(t,X_{t})$, as by Girsanov’s Theorem, $W_{t}=\int_{0}^{t}\lambda^{(k)}(u,X_{u})\,{\rm d}u+W_{t}^{(k)}$ is a ${\mathbb{Q}}[\theta]$-Brownian motion for each $k\in{\mathcal{K}}$, and therefore the ${\mathbb{Q}}[\theta]$-drift of $X$ is equal to $\theta(t,X_{t})$.
To satisfy the condition that ${\mathbb{E}}^{{\mathbb{P}}^{(k)}}\left[\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{P}}^{(k)}}\right]=1$, we impose the Novikov condition.
Assumption 2.2 (Novikov condition).
Let ${\mathcal{A}}$ denote the set of admissible drifts $\theta(\cdot,\cdot)$ satisfying

### v2

@@ Abstract
We consider the problem where an agent aims to combine the views and insights of different experts’ models. Specifically, each expert proposes a diffusion process over a finite time horizon. The agent then combines the experts’ models by minimising the weighted Kullback–Leibler divergence to each of the experts’ models. We show existence and uniqueness of the barycentre model and prove an explicit representation of the Radon–Nikodym derivative relative to the average drift model. We further allow the agent to include their own constraints, resulting in an optimal model that can be seen as a distortion of the experts’ barycentre model to incorporate the agent’s constraints. We propose two deep learning algorithms to approximate the optimal drift of the combined model, allowing for efficient simulations. The first algorithm aims at learning the optimal drift by matching the change of measure, whereas the second algorithm leverages the notion of elicitability to directly estimate the value function. The paper concludes with an extended application to combine implied volatility smile models that were estimated on different datasets.
@@
Mixtures of Experts, Kullback–Leibler, Barycentre, Model Combination, Ensemble Model, Deep learning, Elicitability, Volatility Smiles
@@ 1 Introduction
In many situations a modeller or agent aims to find a single model to, e.g., predict, forecast, classify, or approximate observed patterns. While model selection procedures provide rankings of the different models’ performances — thus allowing the agent to select one model — an alternative to disregarding all but one model is model combination. In the statistical literature, so-called forecast ensemble or superensemble models have been widely applied to weather prediction as they are known to improve deterministic forecasts, see e.g., [14] and reference therein. A variant of a statistical superensemble that allows for predictive distribution functions is to regress a quantity of interest (e.g., air pressure, temperature) onto forecasts stemming from different models [14]. Averaging different predictions has also been used in machine learning, where early examples include bagging predictors [7], Adaptive, Reweighting and Combining (ARCing) algorithms and Adaboost [8]. To improve the architecture of neural networks, [16] propose neural networks ensembles (averaging different neural network architectures), which have been applied in, e.g., credit scoring [26] and oil price forecasting [27]. We also refer to [21] for an application to fusion of generative models.
This work focuses on model combinations of dynamic stochastic models in continuous-time. A related stream of literature, though in discrete-time, is called dynamic time warping (see e.g., [5]), which aims at matching or detecting similarities in time series data by finding a “warping path” that minimises a distance between the time series. Applications of dynamic time warping are widespread and range from classification of genome signals [25], to speech recognition [19], and clustering of financial stocks [12].
@@
In this manuscript, we consider a finite number of experts, each having their own belief about the dynamics of a continuous-time diffusion process. The experts’ models are characterised by different probability measures, under which the process follows the experts’ dynamics. An agent then aims at combining the different experts’ models by finding the probability measure that minimises the weighted Kullback–Leibler (KL) divergence to all models. The resulting dynamics of the stochastic process is described by the barycentre model. Moreover, we allow the agent to include their own views, described by expectations of functions of the terminal values of the process or expected running cost, which results in a constrained barycentre model. Key contributions of this work are as follows. First, we derive the optimal value function of our optimisation problem and a succinct representation of the Radon–Nikodym (RN) derivative of the optimal probability measure to, what we term, the average drift measure. We find that the optimal RN derivative is of Escher type albeit different to the classical solution of maximum entropy, see e.g., [11]. Second, we prove that our constraint optimisation problem is equivalent to an optimisation problem of distorting the expert’s barycentre model to include the agent’s constraints. The latter optimisation problem, that is distorting stochastic processes to include constraints has been studied in [20] and [18]. Third, we propose two deep learning algorithms to approximate the optimal probability measure and thus the dynamics of the constrained barycentre process, allowing for simulation under the constrained barycentre model. Fourth, we apply our framework to combining implied volatility smile models that were estimated on different sets of real data.
@@
Here, we utilise the KL divergence to quantify divergences between probability measures on the path space of stochastic processes. Alternatives could be distances on the space of probability measures stemming from optimal transport theory, such as the Wasserstein distance. There are, however, two caveats: first, calculating a Wasserstein barycentre between random variables is in general an NP hard problem [2], highlighting the curse of dimensionality associated with the Wasserstein distance; second, as we work in continuous time and over a finite time horizon, the Wasserstein distance needs to be replaced with the adapted (also called bicausal) Wasserstein distance. The adapted Wasserstein distance between stochastic processes is only known in special cases [6, 3, 15, 24], and most recent results on barycentre with adapted Wasserstein barycentre pertain to existence and uniqueness [4, 1], while with the KL divergence we obtain a closed-form solution of the optimal RN derivative for arbitrary dimension of the stochastic process.
@@
The manuscript is structured as follows. Section 2 introduces the experts’ models and the agent’s optimisation problem (P). Section 3 establishes the solution to the pure barycentre. In Section 3.1 we consider the special case of two expert models and in Section 3.2 the special case of Ornstein-Uhlenbeck processes. The solution to the KL barycentre with beliefs is detailed in Section 4. Key results are the concise representation of the optimal change of measure (Proposition 4.7) and the recasting of optimisation problem (P) into an optimisation problem of minimally distorting the barycentre measure to include the agent’s constraints (Proposition 4.9). In Section 5, we propose two deep learning algorithms to approximate the drift of the optimally combined model. Section 6.1 illustrates and compares these algorithms on simulated examples. A financial application to combining implied volatility smiles models, that were estimated on different datasets, is provided in Section 6.2.
2 Constrained Kullback–Leibler barycentre
@@
This section first introduces the experts’ models and the optimisation problem — minimising the weighted KL divergence between the experts’ models with constraints — that the agent aims to tackle. Second, using a stochastic control approach, we solve the optimisation problem and characterise the optimal model. Third, we relate the optimisation problem to a modified barycentre model, in the spirit of [18].
@@ 2.1 Experts’ opinions and agent’s problem
We work on a complete and filtered probability space, denoted by $\left(\Omega,\mathbb{P},{\mathcal{F}},({\mathcal{F}}_{t})_{t\in{[0,T]}}\right)$, satisfying the usual conditions of right continuity, on which we have a $d$-dimensional Brownian motion $B=(B_{t})_{t\in{[0,T]}}$ (with independent components) and the filtration is the natural one generated by $B$. We further have a $d$-dimensional process $X=(X_{t})_{t\in{[0,T]}}$, that the agent is interested in modelling, and a set of experts ${\mathcal{K}}:=\{1,\dots,K\}$. Each expert $k\in{\mathcal{K}}$ has a probability measure ${\mathbb{P}}^{(k)}$ — called the “model” of expert $k$ — which is equivalent to ${\mathbb{P}}$ (${\mathbb{P}}\sim{\mathbb{P}}^{(k)}$), and believes that under ${\mathbb{P}}^{(k)}$ the process $X$ satisfies the stochastic differential equation (SDE)
@@
where $W^{(k)}=(W^{(k)}_{t})_{t\in{[0,T]}}$ denotes a $d$-dimensional ${\mathbb{P}}^{(k)}$-Brownian motion, $\mu^{(k)}(t,x):{\mathds{R}}_{+}\times{\mathds{R}}^{d}\to{\mathds{R}}^{d}$ the drift, and $\sigma(t,x):{\mathds{R}}_{+}\times{\mathds{R}}^{d}\to{\rm{S}}^{d}_{++}$, where ${\rm{S}}^{d}_{++}$ is the set of $d$-dimensional matrices such that (s.t.) $\sigma^{\intercal}\sigma$ is strictly positive definite, the volatility. That is, each expert has a different view on the drift of the process $X$. If the experts had differing views on the volatility, so that expert–$k$ wanted a diffusion coefficient of $\sigma^{(k)}(t,X_{t})$, then the KL divergence between the experts’ models is infinite, and hence we could not meaningfully determine a weighted KL barycentre model. Instead, one could regularise the problem, by e.g., introducing a new process $Y=(Y_{t})_{t\in{[0,T]}}$ and a new Brownian motion $B^{Y}$ (independent of all others) that mean-reverts to $\sigma^{(k)}(t,X_{t})$. Then the expert model could be defined as
(2a) ${\rm d}X_{t}=\mu^{(k)}(t,X_{t})\,{\rm d}t+Y_{t}\,{\rm d}W_{t}^{(k)}\,,$
(2b) ${\rm d}Y_{t}=\tfrac{1}{\varepsilon}\big{(}\sigma^{(k)}(t,X_{t})-Y_{t}\big{)}\,{\rm d}t+{\rm d}W_{t}^{Y(k)},$
where $W^{Y(k)}$ is a ${\mathbb{P}}^{(k)}$–Brownian motion independent of the other Brownian motions $W^{(k)}$. As $\varepsilon\downarrow 0$, the process $Y$ is pinned to $\sigma^{(k)}(t,X_{t})$. In this manner, the setting where experts have differing volatilities may be cast into our original setting by increasing the state space dimension by $1$.
Returning now to our original setting, the next set of assumptions guarantee that a strong solution to (1) exists.
@@
The functions $\mu^{(k)}(\cdot,\cdot)$, for all $k\in{\mathcal{K}}$, and $\sigma(\cdot,\cdot)$ satisfy the linear growth and Lipschitz continuity conditions. That is here exists a constant $C_{1}<\infty$, such that for all ${t\in[0,T]}$ and all $x\in{\mathds{R}}^{d}$,
@@
where ||$\cdot$|| ⋅ | denotes the Euclidean norm and |||||$\cdot$||| | ⋅ | | the Frobenius norm. Moreover, there exists a constant $C_{2}<\infty$, such that for all ${t\in[0,T]}$ and all $x,y\in{\mathds{R}}^{d}$,
@@
With a slight abuse of notation, when there is no confusion, we use the notation $\mu^{(k)}_{t}:=\mu^{(k)}(t,X_{t})$ and $\sigma_{t}:=\sigma(t,X_{t})$.
While the experts have differing views — in particular they disagree on the drift of $X$ — an agent wishes to combine the opinions of these experts and assigns each expert a weight $\pi=(\pi_{1},\dots,\pi_{K})$ with $\pi_{k}\in[0,1]$, $k\in{\mathcal{K}}$, and $\sum_{k\in{\mathcal{K}}}\pi_{k}=1$. That is, the agent aims at reflecting the experts’ opinions and incorporating them with their own belief into a combined probability measure ${\mathbb{Q}}$, which we call the combined model. The agent’s beliefs are specified as follows: for functions $g\colon{\mathds{R}}_{+}\times{\mathds{R}}^{d}\to{\mathds{R}}$ and $f\colon{\mathds{R}}^{d}\to{\mathds{R}}$ the agent wishes to ensure that under a probability measure ${\mathbb{Q}}$ their beliefs
${\mathbb{E}}^{\mathbb{Q}}[f(X_{T})]=0\qquad\text{and}\qquad{\mathbb{E}}^{\mathbb{Q}}\left[\textstyle\int_{0}^{T}g(u,X_{u}){\rm d}u\right]=0,$
are satisfied. Here, and in the sequel, we use the notation ${\mathbb{E}}^{{\mathbb{Q}}}[\cdot]$ to denote expectation under the agent’s combined model. If $X_{T}$ is univariate and continuously distributed, the choice $f(x)={\mathds{1}}_{\{x<q\}}-\alpha$, $\alpha\in(0,1)$, for example, corresponds to the agent requiring that the Value-at-Risk (VaR) at level $\alpha$ of the combined model equals to $q$. The weights $\pi$ could be, e.g., obtained by computing the posterior probability that the data the agent is using stems from expert-$k$’s model.
We propose that the agent combines the models by finding the (weighted) Kullback–Leibler (KL) barycentre of the expert models subject to the agent’s constraints. Before presenting the formal optimisation problem, we need some conditions on the set of experts’ probability measures that the agent aims to combine. The following is a standing assumption in the paper.
Assumption 2.2 (Compatibility of Expert Models).
For every $k\in{\mathcal{K}}$,
(3) ${\mathbb{E}}^{{\mathbb{P}}^{(k)}}\left[e^{\frac{1}{2}\int_{0}^{T}|{\bar{\gamma}}^{(k)}_{t}|^{2}{\rm d}t}\right]<\infty,$
where ${\mathbb{E}}^{{\mathbb{P}}^{(k)}}[\cdot]$ denotes expectation under the $k^{\text{th}}$ expert’s model, ${\bar{\gamma}}^{(k)}_{t}:=\sigma_{t}^{-1}(\mu_{t}^{(k)}-{\bar{\mu}}_{t})$, and ${\bar{\mu}}_{t}:=\sum_{k\in{\mathcal{K}}}\pi_{k}\mu_{t}^{(k)}$ — the average drift of the experts’ models.
The average drift of the experts’ models ${\bar{\mu}}_{t}$ characterises the probability measure ${\mathbb{Q}}[{\bar{\mu}}]$, which we term the “average drift measure” and which plays a central role in the exposition. Indeed by Theorem 1 in [22], the conditions in 2.2 imply that, for all $k\in{\mathcal{K}}$, ${\mathbb{E}}^{{\mathbb{P}}^{(k)}}\left[\frac{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}{{\rm d}{\mathbb{P}}^{(k)}}\right]=1$, where the probability measure ${\mathbb{Q}}[{\bar{\mu}}]$ is defined via the Radon-Nikodym derivative
(4) $\frac{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}{{\rm d}{\mathbb{P}}^{(k)}}:=e^{-\frac{1}{2}\int_{0}^{T}|{\bar{\gamma}}_{t}^{(k)}|^{2}{\rm d}t-\int_{0}^{T}{\bar{\gamma}}^{(k)\intercal}{\rm d}W^{(k)}_{t}}.$
Moreover, by Girsanov’s Theorem, ${\overline{W}}:=\left(W_{t}^{(k)}+\int_{0}^{t}{\bar{\gamma}}^{(k)}_{u}{\rm d}u\right)_{{t\in[0,T]}}$ is a ${\mathbb{Q}}[{\bar{\mu}}]$–Brownian motion, for every $k\in{\mathcal{K}}$. We also make use of the average drift function ${\bar{\mu}}(t,x):=\displaystyle\sum_{k\in{\mathcal{K}}}\pi_{k}\,\mu^{(k)}(t,x)$.
We are now in a position to state the agent’s optimisation problem:
@@
denotes the KL divergence of ${\mathbb{Q}}$ relative to ${\mathbb{P}}^{(k)}$ and ${\mathcal{Q}}$ is the set of probability measures parametrised as
@@
${\mathcal{Q}}:=\Bigg{\{}\;{\mathbb{Q}}[\theta]~{}\Big{|}~{}$ $\frac{{\rm d}{\mathbb{Q}}[\theta]}{{\rm d}{\mathbb{Q}}[{\bar{\mu}}]}=\exp\left\{-\tfrac{1}{2}\int_{0}^{T}|{\bar{\lambda}}_{t}|^{2}{\rm d}t-\int_{0}^{T}{\bar{\lambda}}^{\intercal}_{t}\,{\rm d}{\overline{W}}_{t}\right\}\,,$
$\text{where}\quad{\bar{\lambda}}_{t}:=\sigma^{-1}_{t}\big{(}{\bar{\mu}}_{t}-\theta_{t}\big{)}\,,$
[…]